Two charges q1 = -6 µC and q2 = +10 µC are placed 12 cm apart. Another charge, q=-2 μC is placed in the middle. Determine the magnitude and direction of the force acting on q3 by the charge q1 and q2.

1
Expert's answer
2022-01-16T13:11:16-0500

Explanations & Calculations


  • It is needed to use Coulomb's law to the 2 charge combinations described here: "\small (-6\mu C\,,\, -2\mu C)\,\&\,(+10\mu C\, , \, -2\mu C)"

"\qquad\qquad
\begin{aligned}
\small F_1&=\small k\frac{q_1.q_2}{r^2}\\
&=\small (9\times10^9\,Nm^2C^{-2}).\frac{(-6\times10^{-6} C)\times (-2\times 10^{-6}C)}{(0.06\,m)^2}\\
&=\small +30\,N\\

\\

\small F_2&=\small (9\times10^9).\frac{(+10\times10^{-6})\times(-2\times10^{-6})}{(0.06)^2}\\
&=\small -50\,N
\end{aligned}"

  • The signs of the forces are different, meaning that the forces are acting not in the same direction rather opposite to each other.
  • Then the total is "\small F_{net}=30-50=-20\,N".
  • If we consider the arrangement is in a way similar to "\small -6\mu C \to -2\mu C\to+10\mu C", then the positive side is the left.
  • Therefore, the net force on the q3 is to the right "\small -2\mu C \to 10\mu C" .